This is for an exam. I need some help on Solving a quadratic equation with imaginary roots. Find all real or imaginary roots of:
3x^2+4x+4=0


As an example a similar problem: x^2+4x+5=0 had a solution of: x= -2+i or -2-i


Thanks

Answer by stanbon(75887)

3x^2+4x+4=0
Use the quadratic formula, as follows:
x=[-4+sqrt(4^2-4(3)(4)]/2(3) or x=[-4-sqrt(4^2-4(3)(4)]/2(3)
x=[-4+sqrt(-32)]/6 or x=[-4-sqrt(-32)]/6
x=[-2+2isqrt2]/3 or x=[-2-2isqrt2]/3
x=(2/3)(-1+isqrt2] or x=(2/3)[-1-isqrt2]
Cheers,
Stan H.